# One Prompt To Rule Them All: LLMs for Opinion Summary Evaluation

### v1

@@ 6 Results and Analysis
G-Eval vs. Op-I-Prompt vs. Op-Prompts. Table 2 and Table 5 report the summary-level ***Check Appendix C for definition. correlation scores on the SummEval-OP and OpinSummEval dataset. In the case of closed-source models, we observe that our Op-I-GPT-3.5 outperforms or performs comparably to G-Eval-3.5 across all dimensions on both datasets. Specifically, our Op-I-GPT-3.5 outperforms G-Eval-3.5 on all $4$ dimensions for the OpinSummEval dataset, whereas for the SummEval-Op dataset, outperforms on coherence, faithfulness, and aspect coverage, performs comparably on relevance and sentiment consistency, underperforms slightly on fluency  and specificity.
@@
FL $*{}^{*}$ 0.48 $\mathbf{2.9\times 10^{-4}}$ $\mathbf{2.6\times 10^{-4}}$
@@
CO $*{}^{*}$ 0.52 $\mathbf{2.1\times 10^{-4}}$ $\mathbf{3.2\times 10^{-8}}$
@@
RE G-Eval-Mistral 0.51 $6.7\times 10^{-2}$ $\mathbf{4.6\times 10^{-2}}$
@@
$*{}^{*}$ 0.63
Table 4: Significance Test. P-values computed using Mann-Whitney U Test (MW) and T-Test (TT) between the average Spearman correlation scores (AVG-S) taken over $10$ independent generations from G-Eval-Mistral and Op-I-Mistral. Bold for AVG-S indicates better performance, and for MW and TT indicates p-value $<0.05$. *** represents significant performance.
Significance Testing. We perform significance testing using the Mann-Whitney U Test (McKnight and Najab, 2010) for comparison between Op-I-Mistral and G-Eval-Mistral. Table 2 report results for Spearman and Kendall Tau scores computed by using the scoring function with n=100. Op-I-Mistral significantly (p-value $<0.05$) outperforms G-Eval-Mistral on faithfulness, aspect coverage, sentiment consistency, and specificity. Additionally, we group the $100$ generations into $10$ independent groups and compute Spearman correlations for each group. Table 4 reports the Mann-Whitney U Test and T-Test p-values and arrives at a similar observation of Op-I-Mistral significantly outperforming on aforementioned dimensions except faithfulness.
@@
How sensitive is Op-I-Prompt? We test Op-I-Prompt for $3$ definition variations of the aspect coverage dimension. We paraphrase the original definition (Section 4.2) to create $2$ additional versions, ensuring the meaning is preserved (Appendix E). We let the Op-I-Mistral generate n=100 responses to estimate the score using the scoring function (Section 3.2). The Spearman correlations for the $3$ variations are $0.82$ (Table 2), $0.82$, and $0.81$, indicating that Op-I-Prompt is indifferent to the variations of dimensions’ definition.
7 Conclusion
@@ Appendix A Available Benchmark Dataset
OpinSummEval ${}^{\dagger}$ 0.08 0.06 0.11 0.09 0.14 0.11 0.00 0.00
${}^{\dagger}$ 0.13 0.10 0.13 0.11 0.15 0.11 0.04 0.04
${}^{\dagger}$ 0.13 0.10 0.18 0.15 0.18 0.14 0.07 0.05
${}^{\dagger}$ 0.38 0.30 0.20 0.17 0.20 0.16 0.08 0.06
${}^{\dagger}$ 0.42 0.33 0.35 0.29 0.28 0.22 0.41 0.34
${}^{\dagger}$ 0.06 0.05 0.02 0.01 0.07 0.06 0.20 0.17
${}^{\dagger}$ 0.47 0.42 0.28 0.25 0.34 0.30 0.37 0.33
${}^{\dagger}$ 0.41 0.36 0.29 0.26 0.27 0.23 0.38 0.34
Op-I-GPT-3.5 0.51 0.46 0.36 0.32 0.33 0.29 0.43 0.39
G-Eval-Mistral 0.46 0.41 0.41 0.38 0.36 0.32 0.49 0.45
OP-Mistral 0.35 0.33 0.45 0.41 0.34 0.30 0.45 0.41
OP-I-Mistral 0.46 0.41 0.37 0.35 0.38 0.34 0.49 0.45
Table 5: Spearman ($\rho$) and Kendall Tau ($\tau$) correlations at summary-level on $7$ dimensions for the SummEval-OP dataset. For closed-source, Op-I-Prompt performs comparably to G-Eval, whereas for open-source it outperforms alternatives. $\dagger$ represents results as reported in Shen and Wan (2023)
Appendix B Rater Agreement
We hired $3$ raters who have written papers on opinion summarization (1) or are working in the opinion summarization domain (2). These were male Masters’ students aged 21-30. They were provided with detailed guidelines for evaluating summaries on the $7$ dimensions. All $3$ raters received stipends suitable for the tasks. The annotation interface provided raters with the reviews and associated summaries product-wise. Models associated with summaries were not revealed to the raters to remove any bias.
FL $\downarrow$ CO $\downarrow$ RE $\downarrow$ FA $\downarrow$ AC SC $\downarrow$ SP $\downarrow$
@@ Appendix B Rater Agreement
Figure 5: Performance of different models as rated by human annotators. We observe that GPT-$4$ performs the best followed by Solar-$10.7$B and Mistral-$7$B. Self-supervised models perform worse. In general, all the LLMs perform better than human annotated summaries.
Appendix C Opinion Summary Evaluation
$[\mathcal{M}_{b}(s_{i1}),...,\mathcal{M}_{b}(s_{i\mathcal{N}})])$ (2)
Appendix D Prompts
@@
D.1 Op-I-Prompt for Aspect Coverage
@@
D.2 Op-Prompts for Aspect Coverage
@@
D.3 G-Eval for Aspect Coverage
@@ D.4 Summarization Prompt
Appendix E Dimension Definitions

### v2

@@ 6 Results and Analysis
G-Eval vs. Op-I-Prompt vs. Op-Prompts. Table 2 and Table 6 report the summary-level correlation scores on the SummEval-Op and OpinSummEval dataset. In the case of closed-source models, we observe that our Op-I-GPT-3.5 outperforms or performs comparably to G-Eval-3.5 across all dimensions on both datasets. Specifically, our Op-I-GPT-3.5 outperforms G-Eval-3.5 on all $4$ dimensions for the OpinSummEval dataset, whereas for the SummEval-Op dataset, outperforms on coherence, faithfulness, and aspect coverage, performs comparably on relevance and sentiment consistency, underperforms slightly on fluency  and specificity.
@@
FL G-Eval-Mistral $\mathbf{0.48}$ $\mathbf{2.9\times 10^{-4}}$ $\mathbf{2.6\times 10^{-4}}$
@@
CO G-Eval-Mistral $\mathbf{0.52}$ $\mathbf{2.1\times 10^{-4}}$ $\mathbf{3.2\times 10^{-8}}$
@@
RE G-Eval-Mistral $\mathbf{0.51}$ $6.7\times 10^{-2}$ $\mathbf{4.6\times 10^{-2}}$
@@
Op-I-Mistral $\mathbf{0.63}$
Table 5: Significance Test. P-values computed using Mann-Whitney U Test (MW) and T-Test (TT) between the average Spearman correlation scores (AVG-S) taken over $10$ independent generations from G-Eval-Mistral and Op-I-Mistral. Bold for AVG-S indicates better performance, and for MW and TT indicates p-value $<0.05$. * represents significant performance.
Comparative Analysis. Table 3 shows the model responses for G-Eval and Op-I-Prompt using the Mistral-$7$B model. We observe the following: (a) in general G-Eval erroneously assigns a higher score on average compared to the Op-I-Prompt, and (b) Op-I-Prompt ensures that the responses adhere to a specific structure: initially outlining what is addressed and what is absent, followed by an analysis to determine the appropriate score. In contrast, the G-Eval responses, while mentioning various aspects covered, overlook what is missing, resulting in erroneous higher score assignments.
Significance Testing. We perform significance testing using the Mann-Whitney U Test (McKnight and Najab, 2010) for comparison between Op-I-Mistral and G-Eval-Mistral. Table 2 report results for Spearman and Kendall Tau scores computed by using the scoring function with n=100. Op-I-Mistral significantly (p-value $<0.05$) outperforms G-Eval-Mistral on faithfulness, aspect coverage, sentiment consistency, and specificity. Additionally, we group the $100$ generations into $10$ independent groups and compute Spearman correlations for each group. Table 5 reports the Mann-Whitney U Test and T-Test p-values and arrives at a similar observation of Op-I-Mistral significantly outperforming on aforementioned dimensions (except faithfulness).
@@
How sensitive is Op-I-Prompt? We test Op-I-Prompt for $3$ definition variations of the aspect coverage dimension. We paraphrase the original definition (Section 4.2) to create $2$ additional versions, ensuring the meaning is preserved (Appendix D). We let the Op-I-Mistral generate n=100 responses to estimate the score using the scoring function (Section 3.3). The Spearman correlations for the $3$ variations are $0.82$ (Table 2), $0.82$, and $0.81$, indicating that Op-I-Prompt is indifferent to the variations of dimensions’ definition.
7 Conclusion & Future Work
@@
In the future, we plan to investigate LLMs as evaluators for performing large-scale (all reviews) and multi-source opinion summary evaluations.
@@
Acknowledgements
We express our gratitude to the anonymous reviewers for their valuable feedback. We also extend our thanks to the annotators for their diligent and honest efforts.
@@ Appendix B Rater Agreement
OpinSummEval Rouge-1 $0.08$ $0.06$ $0.11$ $0.09$ $0.14$ $0.11$ $0.00$ $0.00$
Rouge-2 $0.13$ $0.10$ $0.13$ $0.11$ $0.15$ $0.11$ $0.04$ $0.04$
Rouge-L $0.13$ $0.10$ $0.18$ $0.15$ $0.18$ $0.14$ $0.07$ $0.05$
BERTScore $0.38$ $0.30$ $0.20$ $0.17$ $0.20$ $0.16$ $0.08$ $0.06$
BARTScore $0.42$ $0.33$ $0.35$ $0.29$ $0.28$ $0.22$ $0.41$ $0.34$
SummaC $0.06$ $0.05$ $0.02$ $0.01$ $0.07$ $0.06$ $0.20$ $0.17$
ChatGPT-3.5 $0.47$ $0.42$ $0.28$ $0.25$ $\underline{0.34}$ $\underline{0.30}$ $0.37$ $0.33$
G-Eval-3.5 $0.41$ $0.36$ $0.29$ $0.26$ $0.27$ $0.23$ $0.38$ $0.34$
Op-I-GPT-3.5 $\mathbf{0.51}$ $\mathbf{0.46}$ $0.36$ $0.32$ $0.33$ $0.29$ $0.43$ $0.39$
G-Eval-Mistral $\underline{0.46}$ $\underline{0.41}$ $\underline{0.41}$ $\underline{0.38}$ $\underline{0.36}$ $\underline{0.32}$ $\mathbf{0.49}$ $\mathbf{0.45}$
OP-Mistral $0.35$ $0.33$ $\mathbf{0.45}$ $\mathbf{0.41}$ $0.34$ $0.30$ $0.45$ $0.41$
OP-I-Mistral $\underline{0.46}$ $\underline{0.41}$ $0.37$ $0.35$ $\mathbf{0.38}$ $\mathbf{0.34}$ $\mathbf{0.49}$ $\mathbf{0.45}$
Table 6: Spearman ($\rho$) and Kendall Tau ($\tau$) correlations at summary-level on $7$ dimensions for the OpinSummEval dataset. For closed-source, Op-I-Prompt performs comparably to G-Eval, whereas for open-source it outperforms alternatives. $\dagger$ represents results as reported in Shen and Wan (2023)
FL $\downarrow$ CO $\downarrow$ RE $\downarrow$ FA $\downarrow$ AC $\downarrow$ SC $\downarrow$ SP $\downarrow$
@@
Figure 5: Performance of different models as rated by human annotators (Round-II). We observe that GPT-$4$ performs the best followed by Solar-$10.7$B and Mistral-$7$B. Self-supervised models perform worse. In general, all the LLMs perform better than human annotated summaries.
Appendix C Prompts
@@
C.1 Op-I-Prompt for Aspect Coverage
@@
C.2 Op-Prompts for Aspect Coverage
@@
C.3 G-Eval for Aspect Coverage
@@ C.4 Summarization Prompt
Appendix D Dimension Definitions